OpenStudy (anonymous):

x^3-5x>(or equal to) 4x^2. solve each inequality for x?

OpenStudy (anonymous):

-1<=x<=0 x>=5

OpenStudy (anonymous):

but how did you arrive at that?

OpenStudy (anonymous):

x(x^2-4x-5)>=0 x(x-5)(x=1)>=0 x=0,5,-1.